# Rethinking Decoders for Transformer-based Semantic Segmentation: A Compression Perspective

### v3

@@ 3.3 Constructing an Ideal Principal Subspace via Self-Attention
where $\boldsymbol{P}^{\prime}=[\boldsymbol{p}^{\prime}_{1},\ldots,\boldsymbol{p}^{\prime}_{M}]\in\mathbb{R}^{D\times M}$ consists of $M$ different bases in $\boldsymbol{P}$, and $\gamma$ is a product of $M(M-1)$ and a constant w.r.t. $\boldsymbol{Z}$. We thus divide the columns of $\boldsymbol{P}$ into $H=C/M$ non-overlapping groups, denoted as $\boldsymbol{P}^{\prime}_{1},\ldots,\boldsymbol{P}^{\prime}_{H}$, and maximize the lower bound for each group, respectively. We thus reformulate the MSSA operator in our case as:
@@ 4.1 Segmentation Performance
Performance comparisons and analysis. From Table 1 we read that all variants of DEPICT outperform Segmenter with significantly fewer parameters and FLOPS on ADE20K validation set. In particular, the best-performing variant, DEPICT-SA based on ViT-Large with an input resolution of 640$\times$640, surpasses Segmenter by 1.1/0.7 mIoU for single/multi-scale inference, while using only 1/14 of the parameters and 1/3 of the FLOPs. We mainly attribute the efficiency of DEPICT to its removal of the FFN block, which plays no role in our interpretations. Such a redundancy has been observed empirically in the field of NLP [32]. but our ablation study in Appendix C shows that naively adopting MSSA while removing the FFN from Segmenter results in worse performance. Moreover, [14] proves that pure attention causes the rank of tokens to decrease rapidly with depth, which is desirable for us to construct an ideal principal subspace. In particular, according to [36], both a finer segmentation granularity and a more performant backbone can significantly boost segmentation performance, partly explaining the inferiority of DEPICT compared to MaskFormer. In Table 2, we find that PICT performs slightly worse than Segmenter. We believe that this is due to the limited size of the dataset; the training set sizes of ADE20K, Cityscapes, and Pascal Context are about 20K, 3K and 5K, respectively. As Table 3 shows, DEPICT requires a sufficient amount of data to demonstrate its superiority.
@@ 4.2 Desirable Properties of DEPICT
Orthogonal properties. According to our interpretations, the parameter matrices of attention are consisted of orthonormal bases, and the extracted class embeddings are fundamentally related to the principal directions, which are also orthonormal. As shown in Figure 4(c), both $\boldsymbol{P}$ and $\boldsymbol{Q}$ are very close to being orthonormal in DEPICT. Although the parameter vectors learned by Segmenter are also nearly orthogonal, their norms are too small. Despite there are semantic similarities among the predefined classes, the class embeddings of Segmenter are excessively related. Moreover, as we expected in Section 3.5 and shown in Figure 5, the class embeddings of different classes nearly lie in a union of orthogonal subspaces, and thus the image embeddings are very likely to satisfy assumption for LDR.
@@
Measuring the coding rate and robustness. On the ViT-L variant of DEPICT-SA with an input resolution of 512$\times$512, we measure the projected coding rate of image embeddings onto the subspaces spanned by $\boldsymbol{P}^{\prime}_{h}$ across layers, as shown in Figure 6. We find that the sign of the learned step size distinguishes two types of subspaces (distinguished by red and blue regions in the figure): one onto which the projected coding rate increases and the other onto which it decreases, which is consistent with our derivations. Furthermore, we measure the robustness of DEPICT under four types of parameter perturbations and find that DEPICT is surprisingly robust whereas Segmenter collapses, as shown in Table 4. We attribute the robustness of DEPICT to its awareness of and modeling for low-dimensional subspaces, which are not significantly altered under the parameter perturbations. Actually, the seminal work [37] has already noticed that it is the subspace, rather than individual units, that contains the semantic information in the high layers of neural networks. Our work demonstrates that this intriguing property can be strengthened by improving model designs, which yields better robustness.
@@ Appendix B Implementation Details
Variant settings. For DEPICT-SA on the ADE20K dataset, we used three MSSA layers and set #heads×$\times$×dim_head to 3$\times$100 across all variants, with the exception of six MSSA layers for the ViT-L variant. For DEPICT-CA, we used three MSSA layers followed by three MSCA layers, and set #heads×$\times$×dim_head to 3$\times$100 in MSSA and #heads×$\times$×dim_head to 3$\times$50 in MSCA, across all variants, with the exception of setting 3$\times$50 in MSSA for the ViT-S variant and six MSSA layers for the ViT-L variant. Only ViT-L variants of DEPICT-SA are evaluated on Cityscapes and Pascal Context, using two and four MSSA layers, and setting #heads×$\times$×dim_head to 4$\times$19 and 4$\times$60, respectively.
@@ Appendix C Additional Experiments
Figure 7: Measuring $R(\boldsymbol{Z})$ and $R(\boldsymbol{Q})$ across layers. All are based on ViT-L with an input resolution of 640$\times$640.
Table 5: Ablation studies. Left: All are based on ViT-S and compared to the naive implementation which adopts MSSA and ISTA, with setting #heads×$\times$×dim_head to 3$\times$50, #layer to 2 and dropout to 0.0. Right: Investigating the impact of #heads and dim_head.
@@
#head×$\times$×dim_head mIoU
@@ D.2 Evaluate the Quality of Solutions for Low-Rank Approximation
We conclude that the better the image embeddings fit a $(C-1)$-dimensional subspace, the more optimal the $k$-means cluster centroids become. In the extreme case where the image embeddings lie perfectly within a $(C-1)$-dimensional subspace, the $k$-means cluster centroids are the optimal solution.
@@ D.3 Understanding self-attention applied to the concatenation of $\boldsymbol{Z}$ and $\boldsymbol{Q}$
$\overline{\boldsymbol{Z}}\cdot\textrm{softmax}\left(\overline{\boldsymbol{Z}}^{\top}\overline{\boldsymbol{Z}}\right)=[\boldsymbol{Z},\boldsymbol{Q}]\cdot\textrm{softmax}\left(\begin{bmatrix}\boldsymbol{Z}^{\top}\boldsymbol{Z}&\boldsymbol{Z}^{\top}\boldsymbol{Q}\\ \boldsymbol{Q}^{\top}\boldsymbol{Z}&\boldsymbol{Q}^{\top}\boldsymbol{Q}\end{bmatrix}\right).$ (83)

### v4

@@ 3.3 Constructing an Ideal Principal Subspace via Self-Attention
where $\boldsymbol{P}^{\prime}=[\boldsymbol{p}^{\prime}_{1},\ldots,\boldsymbol{p}^{\prime}_{M}]\in\mathbb{R}^{D\times M}$ consists of $M$ different bases in $\boldsymbol{P}$, and $\gamma$ is a product of M(M1)M(M-1) and a constant w.r.t. $\boldsymbol{Z}$. We thus divide the columns of $\boldsymbol{P}$ into $H=C/M$ non-overlapping groups, denoted as $\boldsymbol{P}^{\prime}_{1},\ldots,\boldsymbol{P}^{\prime}_{H}$, and maximize the lower bound for each group, respectively. We thus reformulate the MSSA operator in our case as:
@@ 4.1 Segmentation Performance
Performance comparisons and analysis. From Table 1 we read that all variants of DEPICT outperform Segmenter with significantly fewer parameters and FLOPS on ADE20K validation set. In particular, the best-performing variant, DEPICT-SA based on ViT-Large with an input resolution of $\times640$, surpasses Segmenter by 1.1/0.7 mIoU for single/multi-scale inference, while using only 1/14 of the parameters and 1/3 of the FLOPs. We mainly attribute the efficiency of DEPICT to its removal of the FFN block, which plays no role in our interpretations. Such a redundancy has been observed empirically in the field of NLP [32]. but our ablation study in Appendix C shows that naively adopting MSSA while removing the FFN from Segmenter results in worse performance. Moreover, [14] proves that pure attention causes the rank of tokens to decrease rapidly with depth, which is desirable for us to construct an ideal principal subspace. In particular, according to [36], both a finer segmentation granularity and a more performant backbone can significantly boost segmentation performance, partly explaining the inferiority of DEPICT compared to MaskFormer. In Table 2, we find that PICT performs slightly worse than Segmenter. We believe that this is due to the limited size of the dataset; the training set sizes of ADE20K, Cityscapes, and Pascal Context are about 20K, 3K and 5K, respectively. As Table 3 shows, DEPICT requires a sufficient amount of data to demonstrate its superiority.
@@ 4.2 Desirable Properties of DEPICT
Orthogonal properties. According to our interpretations, the parameter matrices of attention are consisted of orthonormal bases, and the extracted class embeddings are fundamentally related to the principal directions, which are also orthonormal. As shown in Figure 4, both $\boldsymbol{P}$ and $\boldsymbol{Q}$ are very close to being orthonormal in DEPICT. Although the parameter vectors learned by Segmenter are also nearly orthogonal, their norms are too small. Despite there are semantic similarities among the predefined classes, the class embeddings of Segmenter are excessively related. Moreover, as we expected in Section 3.5 and shown in Figure 5, the class embeddings of different classes nearly lie in a union of orthogonal subspaces, and thus the image embeddings are very likely to satisfy assumption for LDR.
@@
Measuring the coding rate and robustness. On the ViT-L variant of DEPICT-SA with an input resolution of $\times512$, we measure the projected coding rate of image embeddings onto the subspaces spanned by $\boldsymbol{P}^{\prime}_{h}$ across layers, as shown in Figure 6. We find that the sign of the learned step size distinguishes two types of subspaces (distinguished by red and blue regions in the figure): one onto which the projected coding rate increases and the other onto which it decreases, which is consistent with our derivations. Furthermore, we measure the robustness of DEPICT under four types of parameter perturbations and find that DEPICT is surprisingly robust whereas Segmenter collapses, as shown in Table 4. We attribute the robustness of DEPICT to its awareness of and modeling for low-dimensional subspaces, which are not significantly altered under the parameter perturbations. Actually, the seminal work [37] has already noticed that it is the subspace, rather than individual units, that contains the semantic information in the high layers of neural networks. Our work demonstrates that this intriguing property can be strengthened by improving model designs, which yields better robustness.
@@ Appendix B Implementation Details
[…]
@@ Appendix C Additional Experiments
Figure 7: Measuring $R(\boldsymbol{Z})$ and $R(\boldsymbol{Q})$ across layers. All are based on ViT-L with an input resolution of $\times640$.
[…]
@@
$\timesdim_head$ mIoU
@@ D.2 Evaluate the Quality of Solutions for Low-Rank Approximation
We conclude that the better the image embeddings fit a (C1)(C-1)-dimensional subspace, the more optimal the $k$-means cluster centroids become. In the extreme case where the image embeddings lie perfectly within a (C1)(C-1)-dimensional subspace, the $k$-means cluster centroids are the optimal solution.
@@ D.3 Understanding self-attention applied to the concatenation of $\boldsymbol{Z}$ and $\boldsymbol{Q}$
[…]